• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Glashausser

    Log entry time 16:01:15 on January24,2002

    Entry number 80568

    keyword=Day shift log

    8 AM In controlled access. Techs working on power supply for collimator.
    Detector gas system interlock reads low pressure. No gas flow
    read for VDC chambers. BW says OK to wait for J Segal around 10.

    8:30 Collimator problem is dead power supply. Have to find it and replace it.

    10 Jack has reset the gas system. There was adequate gas everywhere, including
    nitrogen. Not clear why the system shut down.
    Techs still working on collimator system. Network communication lost during
    IOC reboot. Checking with Javier.
    Rusty working on TV readout of JT valve position.
    HRS L magnets have been shut down during this process.
    JP has been working on target controls. See log note.

    10:15 Jack in Hall working on VDC ringing problem while other work going on.

    10:35 JT readout now good, with lots of plastic cutting down intensity.

    11:05 Ed reports that collimator was, as expected, in the open position. (Visible check).

    The spare amplifier board for the collimator doesn't work. Now likely to rob the board from the HRS R. This would mean R could not be powered up for the
    parasite experiment.

    12 Getting oscillations even with threshold at 10V. (Cf Segal entry 80555.) May have
    to do major soldering job (4-6 hrs).

    14 Ed and Mark have set up collimator with old card but with different readout for the collimator only. However, after some work, they decided that there was no caliration for this readout. Since the sieve slit must have an accurate calibration to be useful, this readout won't work. AN (and BR)decided to postpone the optics running until the collimator problem can be fixed, eg, on a maintenance day.

    We had thought the problem in the P4 plane was in the last group of 16 wires, and Jack had taken out that card. After consultation with Michaels, it was decided that
    the oscillations were in the first group of 16. Jack is now working on the first
    16.

    14:45 Angle of HRS L changed to 33.97deg.

    15 Concern that BeO tgt seems off horizontally by about 2 mm. Will check with beam to see if TV has moved.

    15:20 Ramping up magnet.
    Jack says threshold voltage to first 16 flaky--crack somewhere around connector. He will jumper voltage from next card.

    15:40 VDC now seems stable. Jack coming out.

    15:55 Bottom line of shift: Collimator is in open position, as determined by eye by techs. It can't be moved remotely. See log entry by mark now being written.
    VDC: Seems stable, presumably because threshold voltage is now
    getting to card. Has been checked on dataspy for about 15 minutes. Of course,
    no strong guarantee that it will last.
    Target: Original problem about lack of confirmation of a tgt position
    resolved with reboot. Now uncertainty about horizontal position of tgt.
    Magnet ready. Hall ready for beam.